# Drivel-ology: Challenging LLMs with Interpreting Nonsense with Depth

### v1

@@ Abstract
We introduce Drivelology, a unique linguistic phenomenon characterised as “nonsense with depth” – utterances that are syntactically coherent yet pragmatically paradoxical, emotionally loaded, or rhetorically subversive. While such expressions may resemble surface-level nonsense, they encode implicit meaning requiring contextual inference, moral reasoning, or emotional interpretation. We find that current large language models (LLMs), despite excelling at many natural language processing (NLP) tasks, consistently fail to grasp the layered semantics of Drivelological text. To investigate this, we construct a small but diverse benchmark dataset of over 1,200 meticulously curated examples, with select instances in English, Mandarin, Spanish, French, Japanese, and Korean. Annotation was especially challenging: each of the examples required careful expert review to verify that it truly reflected Drivelological characteristics. The process involved multiple rounds of discussion and adjudication to address disagreements, highlighting the subtle and subjective nature of the Drivelology. We evaluate a range of LLMs on classification, generation, and reasoning tasks. Our results reveal clear limitations of LLMs: models often confuse Drivelology with shallow nonsense, produce incoherent justifications, or miss the implied rhetorical function altogether. These findings highlight a deeper representational gap in LLMs’ pragmatic understanding and challenge the assumption that statistical fluency implies cognitive comprehension. We release our dataset and code to facilitate further research in modelling linguistic depth beyond surface-level coherence.
@@
Yang Wang1, Chenghao Xiao2, Chia-Yi Hsiao2, Zi Yan Chang3, Chi-Li Chen3, Tyler Loakman3, Chenghua Lin1, 1The University of Manchester, 2Durham University, 3The University of Sheffield
@@ 1 Introduction
Large language models (LLMs) have achieved impressive success across a wide range of natural language processing tasks, from machine translation and summarisation to commonsense reasoning and dialogue generation Achiam et al. (2023); Qwen Team (2024); Liu et al. (2024a); Guo et al. (2025). These models exhibit high degrees of fluency, contextual awareness, and even emergent reasoning capabilities. However, whether such performance reflects genuine understanding or merely statistical pattern-matching, remains an open and pressing question Rayhan et al. (2023).
@@
The continuous evolution of Internet language as a distinct linguistic style offers a novel and insightful avenue for exploring the depth of understanding in LLMs Ignacio et al. (2024); Mei et al. (2024). Internet language, characterised by its dynamic evolution and cultural embedding, serves as an effective indicator to assess whether models truly grasp deeper semantics or simply rely on superficial pattern recognition. In particular, we introduce the term Drivelology, combining drivel (i.e., nonsense) with -ology (i.e., the study of), which exemplifies this complexity. Drivelology often involves narratives with dual or multiple layers of meaning, employing non-linear structures and ambiguous expressions that challenge LLMs. Unlike purely nonsensical yet grammatically correct sentences such as “Colourless green ideas sleep furiously” Lees (1957), or simplistic tautologies like “either it is or it isn’t”, Drivelology intentionally embeds subtle cultural references, irony, or satire within superficially trivial or absurd narratives. Hence, it differs significantly from typical internet content, such as inspirational quotes or prose, by demanding deeper interpretative engagement from both human readers and LLMs.
@@
Previous studies have explored the difficulty for LLMs to understand humour, sarcasm, and irony Loakman et al. (2023); Romanowski et al. (2025); Zheng et al. (2025). However, Drivelology differs fundamentally from these phenomena by employing more complex, non-linear narratives and deeper ambiguities, making it a uniquely challenging benchmark for assessing LLMs’ semantic comprehension. Examining the capability of LLMs to comprehend Drivelology is critical, as it provides a unique window into their social and semantic comprehension abilities. As a significant facet of human creative expression, Drivelology encapsulates subtle emotions and culturally embedded meanings, offering valuable insights into the complexities of human communication. Understanding such linguistic forms is essential for developing socially intelligent systems Gandhi et al. (2023); Kosinski (2024); Mittelstädt et al. (2024). Moreover, enhancing AI’s grasp of Drivelology can potentially boost creativity in AI applications, improving user experience in content creation tools Hu et al. (2024), and even aiding model safety Matamoros Fernandez et al. (2023) through better detection of contextually ambiguous content.
To this end, we collected and annotated a novel benchmark dataset called DrivelHub for understanding Drivelology from the internet, focusing on narratives that are nonsense with depth. Each sample is annotated with the underlying meaning it reveals. The annotation process was exceptionally rigorous: every candidate example underwent multiple rounds of independent review by expert annotators, followed by in-depth group discussions to resolve ambiguities and disagreements. Given the subtlety and subjectivity of Drivelological meaning, we established detailed annotation guidelines (Figure 6) and required consensus among annotators before finalising each label. This intensive process ensured high-quality, reliable annotations and reflected the complexity inherent in capturing implicit meaning and layered rhetorical intent. We introduce four tasks: (1) Drivelology Detection: A binary classification task to determine whether a given text is non-Drivelology or Drivelology, (2) Drivelology Tagging: A multi-label classification task to assign one or more of the following categories (§3.1) to Drivelology samples: Misdirection, Paradox, Switchbait, Inversion, and Wordplay, (3) Implicit Narrative Writing: This task involves generating an implicit narrative for a given Drivelology sample, and (4) Narrative Selection: A multiple-choice task where the model selects the correct narrative from five options. These tasks collectively encompass various levels of Drivelology understanding, ranging from literal content comprehension to more sophisticated narrative reasoning, thereby providing a comprehensive assessment of Drivelology understanding capabilities. We conducted extensive experiments using the DrivelHub dataset, evaluating both proprietary and open-source LLMs. We release our annotations and code to facilitate future AI research on understanding human creative expression at https://github.com/ExtraOrdinaryLab/drivelology.
@@ 2 Related Work
LLMs Evaluations.  Recent LLMs have demonstrated remarkable performance in following human instructions and performing various downstream tasks through zero-shot prompting Naveed et al. (2023); Liang et al. (2024); Chang et al. (2024); Liu et al. (2024b). Various benchmarks have been proposed to evaluate their performance, primarily focusing on assessing the fundamental capabilities of LLMs Welbl et al. (2017); Wang et al. (2018); Zellers et al. (2019); Sakaguchi et al. (2021); Hendrycks et al. (2021); Suzgun et al. (2023); Zheng et al. (2023); Zhou et al. (2023); Chollet et al. (2024); Jimenez et al. (2024); Yang et al. (2025). However, the ability of large models to perform in-depth social reasoning and accurately understand human contexts remains underexplored Hu and Shu (2023); Feng et al. (2024).
@@
Humour, Irony, and Sarcasm.  Humour, irony, and sarcasm are fundamental elements of human interaction, each requiring a deep understanding of language, context, and social cues Palmer (2003); Filik et al. (2016); Köder and Falkum (2021); Mir and Laskurain-Ibarluzea (2021); Loakman et al. (2023). While computational approaches have addressed these phenomena Yang et al. (2015); Jentzsch and Kersting (2023); Boutsikaris and Polykalas (2024), they often focus on resolving a central contradiction between a literal statement and its context Kreuz and Link (2002); Misra and Arora (2019). Classic sarcasm, for instance, is typically understood through a single cognitive step: inverting a word’s meaning based on a negative premise, as in, “Forgetting assignments and stressing over grades, what a fun semester,” an example from the BIG-Bench Snarks subset Srivastava et al. (2023). We argue that Drivelology presents a more profound challenge, distinguished by two key characteristics: (1) its compositional, multi-layered structure, and (2) its use of pragmatic paradox and ambiguity. For example, the statement “I deeply admire Che Guevara’s anti-capitalist spirit, so I bought all his merchandise” is not a simple semantic inversion. Its critique of performative activism requires synthesising cultural knowledge, making irony just one component of its layered meaning. Furthermore, Drivelology uses pragmatic paradoxes like “I’m good at everything except what I can’t do.” This statement is not explicitly sarcastic nor ironic; its challenge lies in navigating the ambiguity of the speaker’s intent (e.g., humour, self-deprecation). This reliance on compositional meaning and deliberate ambiguity is what sets Drivelology apart.
@@
Distinguishing Drivelology from Deceptive and Nonsensical Language.  To properly situate our work, it is crucial to distinguish Drivelology from related pragmatic concepts. Cappelen and Dever (2019) identify a category termed deep bullshit: utterances defined by an indifference to whether the words make any sense at all, resulting in genuine nonsense. For example, a statement like “Colourless green ideas sleep furiously” qualifies as deep bullshit as it is semantically null. This is distinct from the more widely discussed Frankfurt-style bullshit, which is characterised by an indifference to truth rather than meaning, often deployed to persuade without regard for fact Frankfurt (2005). For instance, a politician might declare they bring a “fresh perspective, unburdened by the stagnant thinking of Washington insiders”, a statement chosen for its persuasive effect, not its accuracy. Drivelology shares a superficial resemblance with deep bullshit, as both can appear nonsensical. However, the two are fundamentally antithetical in their purpose and construction. Whereas deep bullshit arises from a disregard for meaning, Drivelology is meticulously crafted for the sake of conveying a hidden meaning. It is, as we define it, “nonsense with depth.” The surface-level absurdity of a Drivelological text is a deliberate rhetorical framework, designed to guide an audience toward an implicit critique, observation, or emotional payload. Thus, unlike vacuous deep bullshit, Drivelology is rhetorically complex and purposeful. While other forms of bad language, such as lying or misleading, are defined by their deceptive relationship to truth Cappelen and Dever (2019), Drivelology’s defining feature is its purposeful and creative use of apparent nonsense to generate layered semantics. Clarifying these boundaries is essential for developing AI systems that can appreciate subtle human expression. A truly capable model must differentiate between genuinely meaningless utterances (deep bullshit) and the sophisticated, implicit communication of Drivelology, a task that requires moving beyond surface-level coherence to grasp complex rhetorical intent.
@@ 3 The DrivelHub Dataset
Our benchmark dataset, DrivelHub, is designed to evaluate how well LLMs understand Drivelology. Each entry in the dataset includes: (1) a Drivelology sample, (2) the underlying message that the sample aims to convey, and (3) one or more categories describing the main type of the Drivelology sample. These components form the basis for a variety of tasks that assess different aspects of Drivelology comprehension and reasoning. An overview diagram of the multi-stage process for constructing the DrivelHub dataset is presented in Appendix A.1.
3.1 What is Drivelology?
@@
Drivelology refers to a unique style of language that blends humour, ambiguity, and rhetorical complexity to create statements that are intentionally puzzling or nonsensical. Unlike ordinary nonsense or straightforward jokes, Drivelology often relies on layered meanings, unexpected twists, and linguistic playfulness to engage readers in deeper interpretation or amusement. The defining characteristics of Drivelology can be broadly categorised as follows.
@@
Misdirection.  This technique leads the listener down an expected path before a final twist reveals a different, often more literal or absurd, destination. The humour comes from the subversion of the narrative expectation. Example: “Don’t give up on your dream so easily! Keep sleeping!” (The expected path is motivational encouragement; the twist is a literal interpretation of "dream").
@@
Paradox.  This relies on a statement that appears logically self-contradictory but contains a latent, often humorous or profound, truth. The core of the technique is the clash of seemingly incompatible ideas. Example: “I will not forget this favour until I forget it.” (A logically circular statement that humorously asserts the certainty of remembering).
@@
Switchbait. [^1] The term Switchbait is coined in this work to describe a rhetorical device where an initial phrase (the bait) sets up an expectation that is then subverted by a sudden switch in context, cultural reference, or language.  This technique hinges on a specific word or phrase (the “bait”) that has a double meaning. The initial, obvious meaning is suddenly replaced (the “switch”) by a second, surprising one. Example: “British: You’ve got a gun problem. American: Yeah, at least it’s a modern problem.” (The “bait” is the direct criticism of US gun violence. The “switch” is the American’s reply, which reframes the conversation entirely. It pivots from the initial topic to a dark, sarcastic comparison, implicitly mocking the UK’s issues with knife crime as being less “modern.” The humour is generated by this cynical and culturally-specific counter-attack).
@@
Inversion.  This technique takes a well-known phrase, cliché, or social script and flips it on its head. The humour arises from reversing a familiar structure to create a new, often satirical, meaning. Example: “Other than being good-looking, having a great figure, and having money, I have nothing else.” (This inverts the structure of a humble complaint into an arrogant boast).
@@
Wordplay.  This is the general use of linguistic creativity for humorous effect, often by exploiting the sounds or multiple meanings of words. It includes puns, double entendres, and phonetic similarities where the structure isn’t a specific “switch” or “inversion.” Example: “Do you have any raisins? No? How about a date?” (A classic pun playing on two meanings of the word “date”).
@@ 3.2 Drivelology Collection
To ensure a comprehensive evaluation, we prioritised high diversity in our benchmark dataset by selecting a wide range of topics from different languages. Our data was collected from a variety of popular platforms (e.g., Instagram, Threads, TikTok, Facebook, Line, RedNote, Pinterest, Naver, and YouTube). These platforms were chosen strategically as their largest user demographic falls between 25 to 34 years old [^2] According to Statista’s social media demographics data as of 2025., which aligns well with our research focus since Drivelology content predominantly originates from younger generations Sha (2024). We collected all the Drivelology examples by manually and randomly browsing through these platforms, rather than using search terms. It was difficult to apply search term methods to target Drivelology samples because most of the search results were normal posts rather than Drivelology text. For non-Drivelology samples, we curated content from sources such as famous quotes, proverbs, and Ruozhiba (a popular online forum). These non-Drivelology samples are also multilingual, covering English, Mandarin, Spanish, French, Japanese, and Korean. We further categorised non-Drivelology samples into two types: normal sentences (such as meaningful quotes or proverbs) and pure nonsense (text that lacks logical structure or meaning). A significant proportion of the pure nonsense samples were collected from Ruozhiba.

### v2

@@ Abstract
We introduce Drivelology, a unique linguistic phenomenon characterised as “nonsense with depth” – utterances that are syntactically coherent yet pragmatically paradoxical, emotionally loaded, or rhetorically subversive. While such expressions may resemble surface-level nonsense, they encode implicit meaning requiring contextual inference, moral reasoning, or emotional interpretation. We find that current large language models (LLMs), despite excelling at many natural language processing (NLP) tasks, consistently fail to grasp the layered semantics of Drivelological text. To investigate this, we construct a benchmark dataset of over 1,200+ meticulously curated and diverse examples across English, Mandarin, Spanish, French, Japanese, and Korean. Each example underwent careful expert review to verify its Drivelological characteristics, involving multiple rounds of discussion and adjudication to address disagreements. Using this dataset, we evaluate a range of LLMs on classification, generation, and reasoning tasks. Our results reveal clear limitations of LLMs: models often confuse Drivelology with shallow nonsense, produce incoherent justifications, or miss implied rhetorical functions altogether. These findings highlight a deep representational gap in LLMs’ pragmatic understanding and challenge the assumption that statistical fluency implies cognitive comprehension. We release our dataset [^1] https://huggingface.co/datasets/extraordinarylab/drivel-hub and code [^2] https://github.com/ExtraOrdinaryLab/drivelology. to facilitate further research in modelling linguistic depth beyond surface-level coherence.
[…]
latexText page
@@
Yang Wang1, Chenghao Xiao2, Chia-Yi Hsiao2, Zi Yan Chang3, Chi-Li Chen3, Tyler Loakman3, Chenghua Lin1 1The University of Manchester, 2Durham University, 3The University of Sheffield yang.wang-27@postgrad.manchester.ac.uk,    chenghua.lin@manchester.ac.uk
@@ 1 Introduction
Large language models (LLMs) have achieved impressive success across a wide range of natural language processing tasks, from machine translation and summarisation to commonsense reasoning and dialogue generation Tang et al. (2023); Achiam et al. (2023); Qwen Team (2024); Liu et al. (2024a); Guo et al. (2025); Wang et al. (2025); Goldsack et al. (2025). These models exhibit high degrees of fluency, contextual awareness, and even emergent reasoning capabilities. However, whether such performance reflects genuine understanding or merely statistical pattern-matching remains an open and pressing question Bender et al. (2021); Rayhan et al. (2023).
@@
The continuous evolution of Internet language as a distinct linguistic style offers a novel and insightful avenue for exploring the depth of understanding in LLMs Ignacio et al. (2024); Mei et al. (2024). Internet language, characterised by its dynamic evolution and cultural embedding, serves as an effective indicator to assess whether models truly grasp deeper semantics or simply rely on superficial pattern recognition. In particular, we introduce the term Drivelology, combining drivel (i.e., nonsense) with -ology (i.e., the study of), which exemplifies this complexity. Drivelology often involves narratives with dual or multiple layers of meaning, employing non-linear structures and ambiguous expressions that challenge LLMs. Unlike purely nonsensical yet grammatically correct sentences such as “Colourless green ideas sleep furiouslyChomsky (1957), or simplistic tautologies like “either it is or it isn’t”, Drivelology intentionally embeds subtle cultural references, irony, or satire within superficially trivial or absurd narratives. For example, “I deeply admire Che Guevara’s anti-capitalist spirit, so I bought all his merchandise” illustrates how Drivelology paradoxically critiques performative activism. Thus, it differs from typical internet content, such as inspirational quotes or prose, by demanding deeper interpretative engagement from both human readers and LLMs.
@@
Existing studies have explored the difficulty for LLMs to understand humour, sarcasm, and irony Loakman et al. (2023, 2025); Romanowski et al. (2025); Zheng et al. (2025). However, Drivelology differs fundamentally from these phenomena by employing more complex narratives and deeper ambiguities, making it a uniquely challenging benchmark for assessing LLMs’ semantic comprehension. Studying LLMs’ ability to handle Drivelology offers insights into their social and semantic reasoning, as it encodes subtle emotions and culturally embedded meanings. Understanding such linguistic forms is essential for developing socially intelligent systems Gandhi et al. (2023); Kosinski (2024); Mittelstädt et al. (2024). Moreover, enhancing AI’s grasp of Drivelology can potentially boost creativity in AI applications, improving user experience in content creation tools Hu et al. (2024), and even contribute to model safety Matamoros Fernandez et al. (2023) through better understanding of contextually ambiguous content.
Our contributions are as follows:
  • We design a novel taxonomy for Drivelological narratives to aid in categorising the source of meaning embedded in the text.
  • We collect and rigorously annotate a novel benchmark dataset called DrivelHub for understanding Drivelology from the internet, consisting of 1,200+ Drivelological examples that are considered nonsense with depth.
  • We use DrivelHub as the basis for four novel tasks: (1) Drivelology Detection: A binary classification task to determine whether a given text is Drivelology or non-Drivelology; (2) Drivelology Tagging: A multi-label classification task to assign one or more categories from our taxonomy (§3.1) to Drivelology samples; (3) Implicit Narrative Writing: An implicit narrative explanation task for a given Drivelology sample; and (4) Narrative Selection: A multiple-choice task where the model selects the correct narrative from five options.
These tasks collectively encompass various levels of Drivelology understanding, ranging from literal content comprehension to more sophisticated narrative reasoning, thereby providing a comprehensive assessment of Drivelology understanding capabilities. We conducted extensive experiments using the DrivelHub dataset, evaluating both proprietary and open-source LLMs.
@@ 2 Related Work
LLMs Evaluations.  Recent LLMs have demonstrated remarkable performance in following human instructions and performing various downstream tasks through zero-shot prompting Naveed et al. (2023); Liang et al. (2024); Chang et al. (2024); Liu et al. (2024c, b). Various benchmarks have been proposed to evaluate their performance, primarily focusing on assessing the fundamental capabilities of LLMs Zellers et al. (2019); Sakaguchi et al. (2021); Hendrycks et al. (2021); Suzgun et al. (2023); Zheng et al. (2023); Zhou et al. (2023); Jimenez et al. (2024); Yang et al. (2025); Wang and Lin (2025); Hong et al. (2025). However, the ability of large models to perform in-depth social reasoning and accurately understand human contexts remains underexplored Hu and Shu (2023); Feng et al. (2024).
@@
Humour, Irony, and Sarcasm.  Humour, irony, and sarcasm are fundamental elements of human interaction, each requiring a deep understanding of language, context, and social cues Palmer (2003); Filik et al. (2016); Köder and Falkum (2021); Mir and Laskurain-Ibarluzea (2021); Loakman et al. (2023, 2025). While computational approaches have addressed these phenomena Yang et al. (2015); Jentzsch and Kersting (2023); Boutsikaris and Polykalas (2024), they often focus on resolving a central contradiction between a literal statement and its context Kreuz and Link (2002); Misra and Arora (2019). Classic sarcasm, for example, is typically understood through a single cognitive step: inverting a word’s meaning based on a negative premise, as in, “Forgetting assignments and stressing over grades, what a fun semester”.
We argue that Drivelology presents a more profound challenge, distinguished by two key characteristics: (1) its compositional, multi-layered structure, and (2) its use of pragmatic paradox and ambiguity. For example, “I deeply admire Che Guevara’s anti-capitalist spirit, so I bought all his merchandise” is not a simple semantic inversion. Its critique of performative activism requires synthesising cultural knowledge, making irony just one component of its layered meaning. Furthermore, Drivelology uses pragmatic paradoxes like “I’m good at everything except what I can’t do.” This statement is not explicitly sarcastic nor ironic; its challenge lies in navigating the ambiguity of the speaker’s intent. This reliance on compositional meaning and deliberate ambiguity is what sets Drivelology apart.
@@
Distinguishing Drivelology from Deceptive and Nonsensical Language.  To properly situate our work, it is crucial to distinguish Drivelology from related pragmatic concepts. Cappelen and Dever (2019) identify a category termed deep bullshit: utterances defined by an indifference to whether the words make any sense at all, resulting in genuine nonsense. For example, a statement like “Colourless green ideas sleep furiouslyChomsky (1957) qualifies as deep bullshit as it is semantically null. This is distinct from the more widely discussed Frankfurt-style bullshit, which is characterised by an indifference to truth rather than meaning, often deployed to persuade without regard for fact Frankfurt (2005). For instance, a politician might declare they bring a “fresh perspective, unburdened by the stagnant thinking of Washington insiders”, a statement chosen for its persuasive effect, not its accuracy. Drivelology shares a superficial resemblance with deep bullshit, as both can appear nonsensical. However, the two are fundamentally antithetical in their purpose and construction. Whereas deep bullshit arises from a disregard for meaning, Drivelology is meticulously crafted for the sake of conveying a hidden meaning. It is, as we define it, “nonsense with depth”.
The surface-level absurdity of a Drivelological text is a deliberate rhetorical framework, designed to guide an audience toward an implicit critique, observation, or emotional payload. Thus, unlike vacuous deep bullshit, Drivelology is rhetorically complex and purposeful. While other forms of bad language, such as lying or misleading, are defined by their deceptive relationship to truth Cappelen and Dever (2019), Drivelology’s defining feature is its purposeful and creative use of apparent nonsense to generate layered semantics. Clarifying these boundaries is essential for developing AI systems that can appreciate subtle human expression. A truly capable model must differentiate between genuinely meaningless utterances (deep bullshit) and the sophisticated, implicit communication of Drivelology, a task that requires moving beyond surface-level coherence to grasp complex rhetorical intent.
@@ 3 The DrivelHub Dataset
Our benchmark dataset, DrivelHub, is designed to evaluate how well LLMs understand Drivelology. Each entry in the dataset includes: (1) a Drivelology sample, (2) the underlying message that the sample aims to convey, and (3) one or more categories describing the main type of Drivelology contained in the sample. These components form the basis for a variety of tasks that assess different aspects of Drivelology comprehension and reasoning. An overview diagram of the multi-stage process for constructing the DrivelHub dataset is presented in Figure 1 and Appendix A.1.
3.1 Taxonomy of Drivelology
@@
Drivelology refers to a unique style of language that blends humour, ambiguity, and rhetorical complexity to create statements that are intentionally puzzling or nonsensical. Unlike ordinary nonsense or straightforward jokes, Drivelology often relies on layered meanings, unexpected twists, and linguistic playfulness to engage readers in deeper interpretation or amusement. The defining characteristics of Drivelology can be broadly categorised into the following taxonomy:
@@
Misdirection.  This technique leads the listener down an expected path before a final twist reveals a different, often more literal or absurd, ending. Example: “Don’t give up on your dream so easily! Keep sleeping!” The expected path is motivational encouragement; the twist is a literal interpretation of “dream.”
@@
Paradox.  This relies on a statement that appears logically self-contradictory but contains a latent, often humorous or profound truth. The core of the technique is the clash of seemingly incompatible ideas. Example: “I will not forget this favour until I forget it.” This is a logically circular statement that humorously asserts the certainty of remembering.
@@
Switchbait.   This technique hinges on a specific phrase (the “bait”) that has a culturally-embedded double meaning. The initial context is then suddenly replaced (the “switch”) by a surprising second meaning. The humour is generated by this cynical or culturally-specific reinterpretation of the bait, rather than by derailing a narrative. Example: “Brit: You’ve got a gun problem. American: Yeah, at least it’s a modern problem.” The bait is the phrase “gun problem.” The switch reframes it from a criticism of US gun violence to a dark turn, implying cultural counter-attack on UK knife crime.
@@
Inversion.  This technique takes a well-known phrase, cliché, or social script and flips it on its head. The humour arises by reversing a familiar structure to creating a new, often satirical, meaning. Example: “Other than being good-looking, having a great figure, and having money, I have nothing else.” This inverts the structure of a humble complaint into an arrogant boast.
@@
Wordplay.  This is the use of linguistic creativity, often by exploiting the phonetics or polysemy of words. It includes puns, double entendres, and similarities. Example: “Do you have any raisins? No? How about a date?” This is a classic homographic pun playing on two meanings of the word “date”.
@@ 3.2 Drivelology Collection
To ensure a comprehensive evaluation, we prioritised high diversity in our benchmark dataset by selecting a wide range of topics from different languages. Our data was collected from a variety of popular platforms (e.g., Instagram, Threads, TikTok, Facebook, Line, RedNote, Pinterest, Naver, and YouTube). These platforms were chosen strategically as their largest user demographic falls between 25 to 34 years old, [^3] According to Statista’s social media demographics data as of September 2025. https://www.statista.com/topics/1164/social-networks which aligns well with our research focus since Drivelology content predominantly originates from younger generations Sha (2024). For non-Drivelology samples, we curated content from sources such as famous quotes, proverbs, and Ruozhiba (a popular online forum). These non-Drivelology samples are also multilingual, covering English, Mandarin, Spanish, French, Japanese, and Korean. We further categorised non-Drivelology samples into two types: normal sentences (such as meaningful quotes or proverbs) and pure nonsense (text that lacks logical structure or meaning). A significant proportion of the pure nonsense samples were collected from Ruozhiba.